# Difference-in-Discontinuities: Estimation, Inference and Validity Tests

### v1

@@ Abstract
This paper investigates the econometric theory behind the newly developed difference-in-discontinuities design (DiDC). Despite its increasing use in applied research, there are currently limited studies of its properties. The method combines elements of regression discontinuity (RDD) and difference-in-differences (DiD) designs, allowing researchers to eliminate the effects of potential confounders at the discontinuity. We formalize the difference-in-discontinuity theory by stating the identification assumptions and proposing a nonparametric estimator, deriving its asymptotic properties and examining the scenarios in which the DiDC has desirable bias properties when compared to the standard RDD. We also provide comprehensive tests for one of the identification assumption of the DiDC. Monte Carlo simulation studies show that the estimators have good performance in finite samples. Finally, we revisit Grembi et al. (2016), that studies the effects of relaxing fiscal rules on public finance outcomes in Italian municipalities. The results show that the proposed estimator exhibits substantially smaller confidence intervals for the estimated effects.
Keywords: Difference-in-discontinuities, regression discontinuity design, difference-in-difference
@@ 1 Introduction
The difference-in-discontinuities design (DiDC) aims to solve the limitations of both regression discontinuity designs (RDD) and difference-in-difference designs (DiD) by combining both the temporal and discontinuity-based sources of variation from the data-generating processes (DGP). Grembi et al. (2016) and Eggers et al. (2018) proposed this quasi-experimental approach, examining the difference between pre- and post-treatment periods around a threshold for both treated and not treated groups. Although its use in applied microeconomics works is increasing (Azuaga and Sampaio, 2017; Chicoine, 2017; García, 2020; Albright, 2023), research on the econometric theory of the DiDC, including identification assumptions, estimation, and asymptotic properties, remains limited.
@@
The method offers more flexibility over the standard RDD in some settings and can often be used in contexts where neither the RDD nor the DiD is applicable. In particular, the DiDC can handle cases where control and treatment groups differ significantly so as not to satisfy the parallel trends assumption of the DiD, or when multiple time-invariant confounders exist at the threshold in an RDD setting. Additionally, by incorporating more information into the estimation, the DiDC eliminates the bias in the RDD estimates providing more accurate and reliable estimates of treatment effects under certain assumptions, including ones on the DGP’s functional form.
In this paper, we develop the econometric theory for the design with a discontinuity-based approach to the differences. While Galindo-Silva et al. (2021) develop an identification theory for the fuzzy difference-in-discontinuity design based on the difference of RDD estimations, this work will focus on the complementary approach of studying the discontinuity of the differences.
Drawing from standard assumptions in cross-sectional RDD and assumptions specific to the DiDC framework, we establish conditions for reliable identification. These new assumptions, specific to the DiDC design, restrict how the confounding effects behave around the threshold and through time.
We show that the parameter of interest can be recovered using local polynomial estimation of the differences in the outcomes employing the methodology proposed by Calonico et al. (2014). We derive the asymptotic properties of the estimator, highlighting scenarios where the asymptotic bias of the DiDC can be smaller than that of the RDD. Additionally, we introduce straightforward tests to evaluate the testable implications of necessary assumptions for identification. We explore the conditions under which DiDC mitigates bias more effectively than the RDD offering practical guidance for unbiased treatment effect estimation.
Through Monte Carlo simulations, we evaluate the finite-sample properties of the estimator comparing its performance to that of the local linear RDD estimator proposed by Calonico et al. (2014) and the nonparametric DiD regression estimator proposed by Sant’Anna and Zhao (2020). By comparing DiDC’s performance with DiD and RDD methods, we identify the scenarios where DiDC outperforms RDDs and we determine the appropriate application contexts for each method. Finally, we apply our estimator to the study by Grembi et al. (2016) on the impact of fiscal rules on municipal deficits in Italy.
[…]
@@
The advantages of the DiDC have been explored across several applied microeconomic studies. For instance, Azuaga and Sampaio (2017) uses the DiDC to analyze the impact of introducing legislation on domestic violence in Brazil. Butts (2021) explore the use of Diff-in-Disc in geographic settings, Chicoine (2017) investigates the expiration of the Assault Weapon Ban (AWB) by comparing municipalities in which the incumbent mayoral party wins a close election with those where the incumbent is defeated and Albright (2023) isolates the causal effects of algorithmic recommendations on decision-makers using DiDC. These real-life examples demonstrate how effective this approach can be in various research settings, highlighting its value as a practical analysis method.
@@
The remainder of this article is organized as follows. Section 2 presents the DiDC as the discontinuity of differences in a potential outcomes model and shows the main identification results, along with the necessary assumptions. Estimation procedures for the treatment effect in the sharp setting are presented in Section 3, along with the derivation of large-sample properties, optimal bandwidths and robust confidence intervals. In Section 4 we present tests for 2 of the identifying assumptions. Monte Carlo simulations are conducted in Section 5 to examine the properties of the estimators. Section 6 provides an empirical illustration and Section 7 concludes. In the Appendix we provide detailed notation, proofs and other methodological results.
2 Difference-in-Discontinuities as the discontinuity of differences
2.1 Setup
[…]
We define four potential outcomes ($Y_{i,t}(D_{1},D_{0})$) for each unit at each $t$, which depend on the treatment of interest and the confounding treatment:
$Y_{i,t}(0,0)\quad Y_{i,t}(0,1)\quad Y_{i,t}(1,0)\quad Y_{i,t}(1,1)$
We are interested in the local causal effect of the treatment of interest, $D_{1}=1$, at $Z=z_{0}$. To estimate this we compare the difference in outcomes between pre- and post-treatment periods for units near the threshold, both for treated and untreated groups. This comparison can be approached through two methods: the differences of RDDs and the RDD of the differences. While prior studies have predominantly employed the former, this research opts for the latter approach. This choice is because the RDD of the differences approach offers a closer analogy to the well-established and robust RDD literature. However, it is essential to highlight that both methods lead to the same estimand, ultimately identifying the same underlying treatment effects, as demonstrated in Claim 1 shown in the next section.
The observed outcome under the perspective of the difference of RDDs is:
$Y_{i,1}$ $=\left[Y_{i,1}\left(1,1\right)D_{0}+Y_{i,1}\left(0,1\right)\left(1-D_{0}\right)\right]D_{1}+\left[Y_{i,1}\left(1,0\right)D_{0}+Y_{i,1}\left(0,0\right)\left(1-D_{0}\right)\right]\left(1-D_{1}\right)$ (1)
$Y_{i,0}$ $=Y_{i,0}\left(1,0\right)D_{0}+Y_{i,0}\left(0,0\right)\left(1-D_{0}\right)$ (2)
Where $Y_{i,0}$ takes this form because there is no treatment $D_{1}$ at time $t=0$.
Meanwhile, the observed outcome under the approach of the RDD of the differences between outcomes at times $t=0$ and $t=1$ is:
$\Delta Y_{i}$ $=\Delta Y_{i}(0)+\left[\Delta Y_{i}(1)-\Delta Y_{i}(0)\right]D_{1}$ (3)
$\Delta Y_{i}\left(D_{1}\right)$ $=Y_{i,1}\left(D_{1}\right)-Y_{i,0}$
$Y_{i,1}\left(D_{1}\right)$ $=Y_{i,1}\left(1,D_{1}\right)+\left[Y_{i,1}\left(1,D_{1}\right)-Y_{i,1}\left(0,D_{1}\right)\right]D_{0}$
$Y_{i,0}$ $=Y_{i,0}\left(0,0\right)+\left[Y_{i,0}\left(1,0\right)-Y_{i,0}\left(0,0\right)\right]D_{0}$
2.2 Identification
Identification in difference-in-discontinuities design relies on the standard assumptions from the cross-sectional RDD given by Hahn et al. (2001) (Assumptions 1 and 2, adapted to include two time-periods), and assumptions concerning the behavior of confounding effects over time (Assumptions 3 and 4). We state the assumptions here.
@@ 2.2 Identification
All potential outcomes are continuous in $Z=z_{0}$, for any $D_{0},D_{1}\in\{0,1\}$ and $t\in\{0,1\}$:
$\lim_{\varepsilon\rightarrow 0}E[Y_{it}(D_{0},D_{1})|Z_{i}=z_{0}+\varepsilon]=\lim_{\varepsilon\rightarrow 0}E[Y_{it}(D_{0},D_{1})|Z_{i}=z_{0}-\varepsilon]$
Assumption 2 (Discontinuity of treatment probability).
The limits $D_{t}^{+}=\lim_{\epsilon\rightarrow 0}E(D_{it}|Z_{it}=z_{0}+\epsilon)$ and $D_{t}^{-}=\lim_{\epsilon\rightarrow 0}E(D_{it}|Z_{it}=z_{0}-\epsilon)$ exist and $D_{t}^{+}\neq D_{t}^{-}$.
Assumption 2 is a relevance condition, which needs to hold for identification in both sharp and fuzzy settings. In the sharp settings, however, it simplifies to $D_{1}^{+}=1$ and $D_{1}^{-}=0$, while $D_{0}^{+}=D_{0}^{-}=0$.
Claim 1.
Under Assumptions 1 and 2, the difference of RDDs, and the RDD of the differences are the same
$\tau^{DiDC}=\tau_{1}^{RDD}-\tau_{0}^{RDD}=\lim_{\varepsilon\rightarrow 0}E[\Delta Y_{t}|Z_{it}=z_{0}+\epsilon]-\lim_{\varepsilon\rightarrow 0}E[\Delta Y_{t}|Z_{it}=z_{0}-\epsilon]$
@@
To establish the identification of the treatment effect, $E[Y_{i}(1)-Y_{i}(0)|Z_{i}=z_{0}]$ we invoke Assumptions 3 and 4, which are specific to the DiDC design and concern the behavior of the confounding effects around the threshold.
Assumption 3 (Time-invariance of confounding effects).
$E\left[Y_{i,0}(1,0)-Y_{i,0}(0,0)|Z_{i}=z_{0}\right]=E\left[Y_{i,1}(1,0)-Y_{i,1}(0,0)|Z_{i}=z_{0}\right]$
Assumption 3 assumption asserts that the effect of confounding policy $D_{0}$ when there is no treatment remains constant over time. That is, any differences between the treatment and control groups at the threshold in $t=1$, not caused by the treatment, should have existed in $t=0$ before the treatment was introduced. The importance of this assumption cannot be understated. Its validity directly affects the credibility of all subsequent estimates and without it, the estimates from DiDC analyses are unreliable and biased. In Section 4.1, we introduce a comprehensive test to assess its validity.
Assumption 4 (Treatment effect is independent from the confounding policy).
$E\left[Y_{i,t}(1,1)-Y_{i,t}(1,0)|Z_{i}=z_{0}\right]=E\left[Y_{i,t}(0,1)-Y_{i,t}(0,0)|Z_{i}=z_{0}\right]=E\left[Y_{i,t}(1)-Y_{i,t}(0)|Z_{i}=z_{0}\right]$
This final assumption needed for identification states that the treatment effect should not be affected by the confounding policy. This condition implies that the total effect captured by an RDD at time $t=1$ would be the sum of two effects: the effect of the confounding treatment and the effect of the treatment of interest. That is, there is no complementarity or interaction between the treatment of interest and the confounding treatment.
Under Assumptions 1 - 4, we derive the Average Treatment Effect of treatment $D_{1}$:
$\tau^{DiDC}$ $=\lim_{\varepsilon\rightarrow 0}E[\Delta Y_{t}|Z_{it}=z_{0}+\epsilon]-\lim_{\varepsilon\rightarrow 0}E[\Delta Y_{t}|Z_{it}=z_{0}-\epsilon]$
$=E\left[Y_{i}\left(1\right)-Y_{i}\left(0\right)|Z_{i}=z_{0}\right]\left(D_{1}^{+}-D_{1}^{-}\right)$
Therefore the estimand for the DiDC is:
$\hat{\tau}^{DiDC}$ $=\frac{\Delta Y^{+}-\Delta Y^{-}}{\left(D_{1}^{+}-D_{1}^{-}\right)}$ (4)
2.2.1 Relaxing the Independence Assumption
Assumption 4 can be overly restrictive and sometimes there could be scenarios where the confounding policy and the treatment of interest might interact. Rather than assuming additive effects at time $t=1$, we can relax this assumption with one that allows the inclusion of multiplicative effects:

### v2

@@ Abstract
This paper provides a formal econometric framework behind the newly developed difference-in-discontinuities design (DiDC). Despite its increasing use in applied research, there are currently limited studies of its properties. We formalize the theory behind the difference-in-discontinuity approach by stating the identification assumptions, proposing a nonparametric estimator, and deriving its asymptotic properties. We also provide comprehensive tests for one of the identification assumption of the DiDC and sensitivity analysis methods that allow researchers to evaluate the robustness of DiDC estimates under violations of the identifying assumptions. Monte Carlo simulation studies show that the estimators have desirable finite-sample properties. Finally, we revisit Grembi et al. (2016), which studies the effects of relaxing fiscal rules on public finance outcomes. Our results show that most of the qualitative takeaways of the original work are robust to time-varying confounding effects.
Keywords: Difference-in-discontinuities, regression discontinuity design, difference-in-differences
@@ 1 Introduction
The difference-in-discontinuities design (DiDC) aims to address the limitations of both regression discontinuity designs (RDD) and difference-in-differences designs (DiD) by combining temporal and discontinuity-based sources of variation from the data-generating process (DGP). Grembi et al. (2016) and Eggers et al. (2018) proposed this quasi-experimental approach, using differences between the pre- and post-treatment periods around a threshold for both the treated and the untreated groups. Despite its increasing use in applied microeconomics (Azuaga and Sampaio, 2017; Chicoine, 2017; Garcia, 2022; Albright, 2024), the econometric theory of DiDC (identification, estimation, and asymptotic behavior) remains limited. Leventer and Nevo (2025) is one of the few papers that study the identification of regression discontinuity in the case of violation of the continuity of potential outcomes assumption and multiple periods. Their estimator can be an alternative to the DiDC.
@@
The method offers more flexibility over the standard RDD in some settings and can often be used in contexts where neither the RDD nor the DiD is applicable. In particular, the DiDC can handle cases where the control and treatment groups differ significantly and do not satisfy the parallel trends assumption of the DiD, or when multiple time-invariant confounders are at or near the threshold in an RDD setting. Additionally, by incorporating more information into the estimation, the DiDC eliminates bias in RDD estimates, providing more accurate and reliable estimates of treatment effects under certain assumptions.
@@
The advantages of the DiDC have been explored in several applied microeconomic studies. For example, Azuaga and Sampaio (2017) use the DiDC to analyze the impact of introducing legislation on domestic violence in Brazil. Butts (2023) explore the use of Diff-in-Disc in geographic settings, Chicoine (2017) investigates the expiration of the Assault Weapon Ban (AWB) by comparing municipalities in which the incumbent mayoral party wins a close election with those where the incumbent is defeated, and Albright (2024) isolates the causal effects of algorithmic recommendations on decision-makers using DiDC. These real-life examples demonstrate how effective this approach can be in various research settings, highlighting its value as a practical analysis method.
In this paper, we develop the econometric theory for the Difference-in-Discontinuity design. While Galindo-Silva et al. (2021) develop an identification theory for the fuzzy difference-in-discontinuity design based on the difference of RDD estimations, this work will focus on the sharp design, developing identification, inference and validity tests for the DiDC design. Leventer and Nevo (2025) also provide an alternative method to the DiDC. In fact, they use the time to address violations of the RDD’s continuity assumption (e.g., the confounding effect at the threshold). However, they do not specifically study the DiDC estimator and compare it to the traditional RDD.
Drawing from standard assumptions in cross-sectional RDD and assumptions specific to the DiDC framework, we establish conditions for reliable identification. These new assumptions, specific to the DiDC design, constrain how confounding effects behave around the threshold and over time.
We show that the parameter of interest can be recovered using local polynomial estimation of the differences in the outcomes employing the methodology proposed by Calonico et al. (2014b). We derive the asymptotic properties of the DiDC estimator and highlight scenarios in which its asymptotic bias can be smaller than that of the RDD. In addition, we introduce simple tests to assess the testable implications of the identifying assumptions. We explore the conditions under which DiDC mitigates bias more effectively than RDD, offering practical guidance for unbiased treatment-effect estimation. We also provide partial identification results for settings in which the identifying assumptions fail to hold.
Through Monte Carlo simulations, we evaluate the finite-sample properties of the estimator, comparing its performance to that of the local linear RDD estimator proposed by Calonico et al. (2014b) and the nonparametric DiD regression estimator proposed by Sant’Anna and Zhao (2020). By comparing DiDC’s performance with DiD and RDD methods, we identify the scenarios where DiDC outperforms scenarios where DiDC outperforms RDDs and contexts for each method. Finally, we apply our estimator to the study by Grembi et al. (2016) on the impact of fiscal rules on municipal deficits in Italy.
This work is closely connected to Calonico et al. (2014b), whose contributions have been instrumental in robust nonparametric RDD estimation and whose estimation methods are used for our DiDC approach. Frölich and Sperlich (2019) discusses the potential intersection between RDD and DiD, but does not provide a formal derivation of the econometric properties, offering only high-level remarks on identification and estimation. This work also aligns with the broader literature on the intersection of RDD and panel data. Pettersson-Lidbom (2012) explore settings where RDD is combined with fixed effects to address small sample issues and violations of the continuous support assumption. Lemieux and Milligan (2008), use a first-difference RD approach to eliminate individual-specific fixed effects by capitalizing on the longitudinal nature of the Finnish Census data. Lastly, Cellini et al. (2010) introduce "dynamic RD" models, accommodating scenarios with multiple treatment opportunities and examining the dynamics of treatment effects.
@@
The remainder of this article is organized as follows. Section 2 presents the DiDC as the discontinuity of differences in a potential outcomes model and shows the main identification results, along with the necessary assumptions. Estimation procedures for the treatment effect in the sharp setting are presented in Section 3, along with the derivation of large-sample properties, optimal bandwidths and robust confidence intervals. In Section 4 we present tests for 2 of the identifying assumptions and in Section 5 we consider the partial identification of treatment effects when the assumptions described in Section 2 are violated. Monte Carlo simulations are conducted in Section 6 to examine the properties of the estimators. Section 7 provides an empirical illustration and Section 8 concludes. In the Appendix we provide detailed notation, proofs and other methodological results.
2 Identification
2.1 Framework
Our model is best suited for panel data. We consider a setting with two time periods ($t\in\left\{0,1\right\}$) and $n$ units, indexed by $i$. For each unit, we observe an outcome $Y_{i,t}$ at each time period and a running variable $Z_{i}$, which is assumed to be time-invariant.
We are interested in measuring the effect of a binary treatment introduced between time periods 0 and 1, denoted as an indicator function $D_{i,1}$. Treatment assignment is a function of the running variable $Z_{i}$. Units with the running variable above a cutoff $z_{0}$ are assigned to treatment, whereas units with the running variable below $z_{0}$ are assigned to control. That is, the assignment rule for treatment $D_{i,1}$ can be formalized as $D_{i,1}=\mathbf{1}\left\{Z_{i}\geq z_{0},t=1\right\}$.
Our setting differs from the canonical RD setting due to the presence of a confounding policy, denoted by $D_{i,0}$, which is introduced before the treatment of interest, but following the same assignment rule around the cutoff $z_{0}$. The assignment rule for the confounding policy is $D_{i,0}=\mathbf{1}\left\{Z_{i}\geq z_{0},t\geq 0\right\}$.
For concreteness, and in anticipation of the empirical analysis, let $Y_{it}$ denote a fiscal outcome of municipality $i$ in period $t$. The running variable $Z_{i}$ is the population of municipality $i$, the confounding policy $D_{i,0}$ is a high wage for the mayor of municipality $i$, and the treatment of interest $D_{i,1}$ is a relaxation in the fiscal rule for municipality $i$.
Potential outcomes are defined as a function of both the treatment of interest and the confounding policy. We define $Y_{i,t}(d_{0},d_{1})$ as the potential outcome for unit $i$ in period $t$ if the confounder and the treatment are set to $d_{0}$ and $d_{1}$, where $(d_{0},d_{1})\in\left\{0,1\right\}^{2}$.
The observed outcomes at each time period are related to the potential outcomes through
$Y_{i,1}$ $=\left[Y_{i,1}\left(1,1\right)D_{i,0}+Y_{i,1}\left(0,1\right)\left(1-D_{i,0}\right)\right]D_{i,1}+$ (1)
$\quad\quad\left[Y_{i,1}\left(1,0\right)D_{i,0}+Y_{i,1}\left(0,0\right)\left(1-D_{i,0}\right)\right]\left(1-D_{i,1}\right)$
and
$Y_{i,0}=Y_{i,0}\left(1,0\right)D_{i,0}+Y_{i,0}\left(0,0\right)\left(1-D_{i,0}\right)$ (2)
Since the policy of interest is only introduced between periods 0 and 1, at period 0 we only observe two of the four possible potential outcomes, whereas in period 1 we can observe all four possible potential outcomes.
The presence of the confounding policy at the threshold is a violation of the continuity of mean potential outcomes assumption invoked in the canonical RD setting (Assumption 1 from Hahn et al. (2001)). In the DiDC setting, the continuity assumption is modified to accommodate the confounding policy:
@@ 2.1 Framework
All potential outcomes are continuous in $Z=z_{0}$. For any $\left(d_{0},d_{1}\right)\in\{0,1\}^{2}$ and $t\in\{0,1\}$:
$\lim_{\varepsilon\rightarrow 0}\mathbb{E}[Y_{it}(d_{0},d_{1})|Z_{i}=z_{0}+\varepsilon]=\lim_{\varepsilon\rightarrow 0}\mathbb{E}[Y_{it}(d_{0},d_{1})|Z_{i}=z_{0}-\varepsilon]$
Assumption 1 is a straightforward modification of the standard continuity assumption. It can be interpreted as stating that, other than the confounder $D_{0}$, there is no systematic difference in mean potential outcomes around the threshold $z_{0}$. In the empirical application, Assumption 1 states that, aside from the wages of mayors and the relaxation of fiscal rules, there are no other policies that change around the population threshold $z_{0}$.
The second identification assumption is the random treatment assignment around the cut-off that is known as non-manipulation at the thresold:
Assumption 2 (Random Treatment Assignment at the Cutoff).
: for $t\in\{0,1\}$
$Y_{i,t}(d_{0},d_{1})\;\perp(D_{i,0},D_{i,1})\mid Z=z_{0}$ (3)
A fundamental assumption in the canonical RD setting is that the probability of receiving treatment is discontinuous at the threshold $z_{0}$. In the DiDC, this assumption is extended for the treatment of interest and the confounding policy:
Assumption 3 (Sharp Discontinuites).
Define the limits $D_{0}^{+}=\lim_{\epsilon\rightarrow 0}\mathbb{E}[D_{i0}|Z_{i}=z_{0}+\epsilon)$, $D_{0}^{-}=\lim_{\epsilon\rightarrow 0}\mathbb{E}[D_{i0}|Z_{i}=z_{0}-\epsilon)$, $D_{1}^{+}=\lim_{\epsilon\rightarrow 0}\mathbb{E}[D_{i1}|Z_{i}=z_{0}+\epsilon)$ and $D_{1}^{-}=\lim_{\epsilon\rightarrow 0}\mathbb{E}[D_{i1}|Z_{i}=z_{0}-\epsilon)$. Assume $D_{i,0}^{+}=D_{i,1}^{+}=1$ and $D_{i,0}^{-}=D_{i,1}^{-}=0$.
Assumption 3 states that there is perfect compliance towards the assignment rule around the threshold $z_{0}$. Borrowing the jargon from the standard RD setting, we call the case with perfect compliance the sharp discontinuity setting. In the sharp setting, we observe $Y_{i,0}=Y_{i,0}(1,0)$ and $Y_{i,1}=Y_{i,1}(1,1)$ for units above the threshold $z_{0}$, whereas we observe $Y_{i,0}=Y_{i,0}(0,0)$ and $Y_{i,1}=Y_{i,1}(0,0)$ for units below the threshold.
Assumptions 1 and 3 modify the traditional RDD assumptions stated at Hahn et al. (2001) to accommodate the confounding policy. However, to identify the effect of the treatment of interest, an additional assumption regarding the evolution of the confounder effect is required:
Assumption 4 (Time-invariance of confounding effects).
$\mathbb{E}\left[Y_{i,0}(1,0)-Y_{i,0}(0,0)|Z_{i}=z_{0}\right]=\mathbb{E}\left[Y_{i,1}(1,0)-Y_{i,1}(0,0)|Z_{i}=z_{0}\right]$
Assumption 4 states that the effect of confounding policy $D_{i,0}$ remains constant over time. That is, any differences between the treatment and control groups at the threshold in $t=1$, not caused by the treatment, should have existed in $t=0$ before the treatment was introduced.
Assumptions 1-4 are the assumptions invoked by Grembi et al. (2016) in order to derive a causal interpretation for the DiDC estimand. In the next section, we discuss the Difference-in-Discontinuities estimand, as well as the Discontinuity-in-Differences estimand.
2.2 The DiDC Estimand
Before discussing identification, we define the relevant target parameters for the setting. There are two relevant causal parameters for understanding the effect of the policy of interest. We are interested in the identification of the average difference between potential outcomes of treated units at the cutoff and the potential outcomes of untreated units, holding the confounder fixed.
Thus, the first parameter we define is $\tau_{c}\equiv\mathbb{E}\left[Y_{i,1}(1,1)-Y_{i,1}(1,0)|Z_{i}=z_{0}\right]$, which is the average treatment effect for individuals at the cutoff exposed to the confounder in period 1. The second parameter of interest is the average treatment effect for individuals at the cutoff not exposed to the confounder in period 1, defined as $\tau_{uc}\equiv\mathbb{E}\left[Y_{i,1}(0,1)-Y_{i,1}(0,0)|Z_{i}=z_{0}\right]$. We now discuss the point identification of these parameters.
We define $\tau_{t}^{RD}$ as the regression discontinuity estimand at period $t$. Let $Y_{t}^{+}=\lim_{\varepsilon\rightarrow 0}E[Y_{i,t}|Z_{i}=z_{0}+\epsilon]$ and $Y_{t}^{-}=\lim_{\varepsilon\rightarrow 0}E[Y_{i,t}|Z_{i}=z_{0}-\epsilon]$. Thus, the RD estimand at period $t$ is simply $\tau_{t}^{RD}=Y_{t}^{+}-Y_{t}^{-}$. Grembi et al. (2016) define the DiDC estimand as the difference between the RD estimand in period 1 and the RD estimand in period 0: $\tau^{DiDC}=\tau_{1}^{RD}-\tau_{0}^{RD}$.
The next lemma shows that the DiDC estimand can be alternatively defined as an RD estimand evaluated at the difference between outcomes over time:
Lemma 1.
Under Assumptions 1, 2 and 3, the difference of RDs, and the RD of the differences are equivalent:
$\tau^{DiDC}=\tau_{1}^{RD}-\tau_{0}^{RD}=\lim_{\varepsilon\rightarrow 0}\mathbb{E}[\Delta Y_{i}|Z_{i}=z_{0}+\epsilon]-\lim_{\varepsilon\rightarrow 0}\mathbb{E}[\Delta Y_{i}|Z_{i}=z_{0}-\epsilon]$
@@ 2.2 The DiDC Estimand
The result in Lemma 1 is intuitive and presents an attractive feature for the DiDC setting: the estimand can be implemented via a single RD estimand rather than taking the difference across estimands. We now turn to the causal interpretation of the estimand, which is already well established in the literature:
Lemma 2 (Grembi et al. (2016)).
Under Assumptions 1-4, we have
$\tau^{DiDC}=\tau_{c}$
The difference-in-discontinuities estimand identifies the effect of the treatment of interest in period 1 for units at the threshold that are exposed to the confounding policy. To identify a more general causal effect, an additional assumption is required:
Assumption 5 (No-interaction between treatment and confounding effects).
$\tau_{c}=\tau_{uc}=\mathbb{E}\left[Y_{i,1}(d_{0},1)-Y_{i,1}(d_{0},0)|Z_{i}=z_{0}\right]$
Assumption 5 states that the effect of the treatment of interest does not depend on the confounding policy. Such an assumption can be justified by a potential outcomes model in which the treatment effect of interest and the effect of the confounder are linearly separable. Assumption 5 might be overly restrictive, as one might expect the treatment of interest and the confounding policy to interact; nevertheless, it allows the DiDC estimand to identify a more general causal effect:
Corollary 1.
Under Assumptions 1-5,
$\tau^{DiDC}=\mathbb{E}\left[Y_{i,1}(d_{0},1)-Y_{i,1}(d_{0},0)|Z_{i}=z_{0}\right]$
2.2.1 Relaxing the No-Interaction Assumption
Assumption 5 can be overly restrictive, and sometimes there could be scenarios where the confounding policy and the treatment of interest might interact. Rather than assuming additive effects at time $t=1$, we can relax this assumption with one that allows the inclusion of multiplicative effects: